# Accreditation Against Limited Adversarial Noise

### v1

@@ Abstract
I present an easily-applicable method for upgrading any pre-existing digital accreditation (a variety of quantum verification) protocol - that assumes IID CPTP error - to function correctly when error is modelled as being adversarial - albeit slightly modified to more closely resemble physical reality - with no diminution in their efficiency or suitability for near-term usage.
@@ I Introduction
Accreditation has provided an efficient, scalable, and practical method for quantifying the the quality of particular computations. This is a vital requirement for making use of quantum computers in the NISQ era [1], when quantum computations will be unreliable due to interactions with the surrounding environment - known as noise [2] - that induce erroneous - i.e. unintended / not in the circuit being executed - operators in a computation; so some notion of whether the outputs of quantum computations are usable - for whatever purpose they are required for - is needed.
I note that accreditation is not the only method of evaluating the quality of quantum computations and/or quantum computers. One such class of protocols is randomized benchmarking [3, 4, 5, 6]. These protocols do not aim to measure the how well a quantum computer can implement a specific computation - e.g. preparing a six-qubit Greenberger-Horne-Zeilinger (GHZ) state [7] - but rather how well it performs benchmarking computations - which usually comprise of a series of gates that, if implemented with no error, return a known measurement outcome with certainty - and hence obtain a measure of the quality of the quantum computer itself, instead of a measure of the quality of a computation’s outputs like accreditation does.
A more useful measure, in the long term, is how well specific computations are implemented, and hence to what extent data obtained from quantum devices can be trusted. Verification [8, 9, 10, 11] is a large field with many protocols aimed at solving exactly this problem. These protocols typically run a quantum computation, allowing for the possibility of error, and - via some machinations - decide whether to accept or reject the returned outputs. This is a more limited measure of the quality of a specific computation than the ideal-actual variation distance (as defined in Def. 1) - which accreditation protocols provide.
@@
Furthermore, verification protocols, broadly speaking, rely on a different kind of assumption to accreditation. Verification has typically used the cryptographic framework of delegated computation  [^1] This is also known as an interactive prover system, wherein successful verification protocols have been shown to exist for a very large set of problems [12, 13], although - ironically - noise may severely restrict when interactive prover systems exist [14].  [15] (i.e. is based on Alice and Bob interacting and precluding error typically [^2] There exists verification protocols that do not have this assumption, that use two quantum devices [16, 17, 18], but here I focus on single-device protocols. either in state preparation [19, 20, 21] or measurement [22, 23, 17, 24]); accreditation has instead tended to base its assumptions on the physics within quantum computers. By turning from the cryptographic assumptions of the verification protocols [8], from which it originates, to physics-based assumptions; accreditation has been able to develop assumptions, based on experimental evidence, and then protocols [25], using the aforementioned assumptions, to initiate a vein of research that has, thus far: provided the aforementioned scalable protocol to achieve confidence in computation outputs, proven itself experimentally implementable and loosened its assumptions [26], and directly led to the first method for quantifying the correctness of the outputs of quantum analogue simulation [27].
@@
But these successes have come at a cost. Most relevantly, for the purposes of this paper, accreditation has hitherto required abandoning the adversarial noise model present in many verification protocols [8]; moving, instead, to a model where noise induces identically and independently distributed (IID) CPTP error [^3] Error which may be represented by a completely positive trace preserving map acting on both the system and its environment.. Returning to an adversarial error model has proven difficult as the physics-based assumptions clash with the adversarial model of noise. This is resolved herein with a “hybrid” approach, using a limited form of adversarial noise wherein the cryptographic setting / Alice and Bob formalism [15], as used in adversarial noise models, is used but modified with a protocol (Protocol 1) for how quantum circuits are executed that both Alice and Bob participate in - still leaving Bob to be as malicious as he likes but limiting, based on the aforementioned physics-based assumptions, his knowledge and abilities. This permits me to improve upon the IID CPTP error model, while retaining validity in all situations where the IID assumption was valid. The result is protocols with less stringent assumptions, loosening the requirements on the physical devices used to implement them.
@@
This is achieved through exploiting limits placed on Bob (based on experimental realities - that the probability of error varies little in executions of the same circuit, on the same hardware, in the same environment - and inspired by single-qubit gates experiencing gate-independent error in pre-existing accreditation protocols [25, 26]) allowing methods to overcome the factors that prohibited adversarial noise in extant accreditation protocols.
@@
This paper proceeds, from here, with an introduction to the error model used - in Sec. II.1 - and its justification - in Sec. II.2. I then present my main result: the adaptation of accreditation protocols to the newly-developed error model in Sec. III. The paper then concludes - in Sec. IV - with a discussion of further possible developments.
II Adversarial Error Model
II.1 Error Model Presentation
@@
My adversarial error model is an adaption of the cryptographic setting [25]. In the cryptographic setting, Alice is attempting to get Bob to perform a sampBQP [28, 29, 30] quantum computation [^4] Meaning a computation to obtain a sample from a specified distribution, that can be efficiently performed on a quantum computer. Note that sampBQP includes both BQP and fBQP (the set of functions efficiently computable on a quantum computer). and provide her with the true results, while only having the ability to perform polynomially-bounded classical computation herself. Bob is computationally unbounded so, if he wished, could implement the computation Alice requires but his aim is instead to trick Alice by providing her with the results of an incorrect computation and to have her believe he has acted honestly i.e. that the results she receives are the outputs of the computation she wanted to be performed when they are not.
@@
My specific error model is slightly modified from the traditional error model used in the cryptographic setting, mainly through limitations on Bob. These limitations are entirely contained in how the circuits Alice requests be executed are executed, which is presented in Protocol 1, but this first requires Def. 2.
@@
A gate in a circuit is said to be redacted if the gate’s position in the circuit - i.e. when it is applied and to which qubits - is specified but the operator it represents - e.g. if it is a Pauli $\hat{X}$ gate or Pauli $\hat{Z}$ gate - is not.
A final concept required to be formally defined before the presentation of how the circuits Alice requests be executed are executed is a notion of the unique choice of CPTP maps representing error in an execution.
@@
A CPTP set corresponding to a specific circuit is a possible set of the CPTP maps that represent the error occurring in the execution of that circuit.
@@
I can then present the protocol by which sets of circuits, under my error model, are executed, in Protocol 1.
@@
    Alice provides the set of circuits (differing only in their single-qubit gates), $\mathcal{S}$, to be executed.
@@
    Bob receives the full details of $\mathcal{S}$, with all single-qubit gates redacted, and picks a set of CPTP sets (each element of which is a valid CPTP set for any element of $\mathcal{S}$) - denoted $\Xi$ subject to the condition that, for all CPTP sets in $\Xi$, the resulting probability of error in the execution of any circuit in $\mathcal{S}$, with any element of $\Xi$ representing its error, is within the interval:
    $\big{[}P_{0}\big{(}1-\beta\big{)},P_{0}\big{(}1+\beta\big{)}\big{]},$ (1)
    for a known (to Alice and Bob, in advance) $\beta\in\mathbb{R}^{+}$.
@@
    For each circuit in $\mathcal{S}$:
@@
      Bob chooses a CPTP set, $\xi$, from $\Xi$.
@@
      The circuit is executed with $\xi$ determining the error occurring during the execution.
@@
      Alice and Bob are both given the measurement outcomes of the execution.
@@
Alice is free to initiate Protocol 1 whenever and as often as she likes; Bob is free to act how he likes to achieve his stated malicious aims but must execute Protocol 1 as prescribed - but is free to act nefariously within the permitted bounds - when it is invoked.
II.2 Physical Origins of the Error Model
The two main limitations on the adversarial error / Bob are:
@@ II.2 Physical Origins of the Error Model
    For any set of circuit executions within a single use of Protocol 1, the respective probabilities of error occurring are all contained in the interval $[P_{0}(1-\beta),P_{0}(1+\beta)]$ (where $\beta\in\mathbb{R}^{+}$ is known to Alice, but $P_{0}$ is not) i.e. that the probability of error is only slightly different in different executions of similar circuits in quick succession. This is represented in Protocol 1 by Eqn. 1.
The first limitation corresponds to single-qubit gates experiencing gate-independent error. This is a standard assumption the the pre-existing accreditation protocols (in Refs. [25, 26, 27]) and follows from single-qubit gates typically being the least error-prone components of a quantum computer [31, 32] (which has held true over time [33, 34], and can be most clearly seen in Ref [35, Fig.5]): the error is so small that the error in different single-qubit gates does not differ much. This is a “standard [assumption] in the literature on noise characterisation and mitigation” [36] and has seen extensive use in theoretical work [3, 37, 38, 39, 40, 41, 42, 43, 44, 45, 36].
@@
    Ref. [26, Figure. 5] ran the accreditation protocol - developed therein - many times, producing a probability of error in the traps for each use of the protocol. The set of error probabilities generated by this can be seen to vary little across many uses of the protocol.
@@
    Ref. [46, Fig. 6, Fig. 7, and Fig. 8] investigated the error rates in single-qubit and two-qubit gates. It found these error rates are very rarely far from their average. Note that the variation shown in this paper is over a much longer time-span (days) than this limitation requires (seconds).
@@
    Ref. [35, Fig. 9] examined NISQ devices and - across many devices - plotted the error rates of different qubits in each device. Fig. 9 shows the error bars on the error rate to be - with some notable exceptions - small fractions of the error rate and varying much more across qubits than for a fixed qubit over time.
@@
    Ref. [47, Fig. 1(b)] looked at the error rates of two-qubit gates on differing pairs of qubits, it shows that - again, over a time-span of days - for a specific pair of qubits, the days when the error rate deviates far from its average are rare, with more extreme deviations being rarer.
Finally, this limitation can be seen as an aim of quantum computer hardware engineering: as quantum computers improve and their actual outputs approach the errorless outputs, with decreasing variance, the typical difference - by any measure - between two executions of similar circuits will tend to zero.
To my knowledge, I am the first to explicitly state and use this second limitation. However, I note that it is already implicitly accepted in the community by the acceptance of randomized benchmarking [3, 4, 5, 6] as a meaningful measure: if even the same circuit repeated multiple times, in very quick succession, produces wildly varying probabilities of error then the variation of error probabilities implies that previous measures mean almost nothing for future computations or the quality of a quantum device over a meaningful time-span. This is not the case, and Ref. [47, Fig. 6] shows that, over a span of days, the results of randomized benchmarking do not vary over a very large range.
There is, however, another, subtle, restriction on Bob: he controls the CPTP maps acting on all executions, but does not control any stochastic elements of the implementation of the CPTP map. I.e. he may specify that a specific CPTP map in the execution applies stochastic Pauli error, but he does not decide if this results in a Pauli $\hat{X}$ gate being applied. Furthermore, Bob is not informed of the outcomes of these stochastic processes - in spite of choosing which stochastic processes occur. This represents that the environment - where any memory of past error that the noise represented by Bob has must be - is ironically not, consistently or over long time-spans, protected from noise acting on it and destroying any information/memory stored in it, so any hostile intelligence that lurks in the noise must have a comparatively short and small memory  [^5] Quantum hardware engineers have expended much effort to keep the the information inside quantum computers intact, so there is no reason to assume the environment can - with non-negligible probability - store significant amounts of information without extensive engineering effort to enable it to do so.. This argument is strengthened by the potential use of error suppression techniques such as dynamic decoupling [48].

### v2

@@ Abstract
I present an accreditation (a variety of quantum verification) protocol where error is assumed to be adversarial (in contrast to the IID assumption used in previous accreditation protocols) - albeit slightly modified to more closely resemble physical reality - with no diminution in efficiency or suitability for near-term usage.
@@ I Introduction
Accreditation provides an efficient and scalable method for quantifying the the quality of quantum computations, without trusting any aspect of the computations. This is a vital requirement for using quantum computers in the NISQ era [1], when quantum computations will be unreliable due to interactions with the surrounding environment - known as noise [2] - that induce erroneous operators in a computation; so some measure of whether the outputs of quantum computations are usable is needed.
I note that accreditation is not the only method of evaluating quantum computations or quantum computers. One such class of protocols is randomized benchmarking [3, 4, 5, 6]. These protocols do not aim to measure the how well a quantum computer can implement a specific computation - e.g. preparing a six-qubit Greenberger-Horne-Zeilinger (GHZ) state [7] - but rather how well it performs benchmarking computations - which usually comprise of a series of gates that, if implemented with no error, return a known measurement outcome with certainty - and hence obtain a measure of the quality of the quantum computer itself, instead of a measure of the quality of a computation’s outputs like accreditation does.
A more useful measure, in the long term, is how well specific computations are implemented and if data obtained from quantum devices can be trusted. Verification [8, 9, 10, 11] is a large field with many protocols aimed at solving exactly this problem. These protocols typically run a quantum computation, allowing for the possibility of error, and - via some machinations - decide whether to accept or reject the returned outputs. This is a more limited measure of the quality of a specific computation than the ideal-actual variation distance (as in Def. 1) which accreditation protocols provide.
@@
Furthermore, verification protocols, broadly speaking, rely on a different kind of assumption to accreditation. Verification has typically used the cryptographic framework of delegated computation [12] (i.e. is based on Alice and Bob interacting and precluding error typically [^1] There exists verification protocols that do not assume there is no error in some particular aspect of a computation, using two quantum devices [13, 14, 15], but here I focus on single-device protocols. either in state preparation [16, 17, 18] or measurement [19, 20, 14, 21]); accreditation has instead tended to base its assumptions on the physics within quantum computers. By turning from the cryptographic assumptions of verification protocols [8], from which it originates, to physics-based assumptions; accreditation has been able to develop assumptions, and then protocols [22] - using those assumptions - to initiate a vein of research that has, thus far: provided the aforementioned scalable protocol to achieve confidence in computation outputs, proven itself experimentally implementable [23], and led to the first method for quantifying the quality of the outputs of quantum analogue simulation [24].
@@
But these successes have come at a cost. Most relevantly, for this paper, accreditation has hitherto required abandoning the adversarial noise model present in many verification protocols [8]; moving, instead, to a model where noise induces identically and independently distributed (IID) CPTP error [^2] Meaning in each execution of a circuit, the error is represented by identical and independent - across multiple executions - completely positive trace preserving maps acting on both the system and its environment.. Returning to an adversarial error model / problem setting has proven difficult as the physics-based assumptions of accreditation protocols clash with the adversarial model of noise. This is resolved herein by upgrading the protocol in Ref. [23] to use/assume a limited form of adversarial noise wherein the cryptographic setting / Alice and Bob formalism [12], as used in adversarial noise models, is used but modified with a protocol (Protocol 1) for how quantum circuits are executed that both Alice and Bob participate in - still leaving Bob to be as malicious as he likes but limiting, based on the aforementioned physics-based assumptions, his knowledge and abilities. This permits me to improve upon the IID CPTP error model, in Ref. [23], while retaining validity in all situations where it was valid. The result is a protocol with less stringent assumptions, loosening the requirements on the physical devices used to implement them.
[…]
Error modelled as CPTP maps Error modelled as CPTP maps
Error is IID Error is adversarial (but limited)
GI single-qubit gate error GI single-qubit gate error
Table 1: Table showcasing the upgrade presented in this paper as a comparison between the assumptions in Ref. [23] and the protocol herein. Note that GI is shorthand for Gate-Independent (see Sec. II.4).
@@
This is achieved through exploiting limits placed on Bob (based on experimental realities - that the probability of error varies little in executions of the same circuit, on the same hardware, in the same environment - and inspired by single-qubit gates experiencing gate-independent (GI) error in pre-existing accreditation protocols [22, 23]) allowing methods to overcome the factors that prohibited adversarial noise in extant accreditation protocols.
@@
This paper proceeds, from here, with an introduction to the problem setting - in Sec. II.3 - and its justification - in Sec. II.4. I then present my main result: the adaptation of accreditation protocols to the newly-developed problem setting / model of error in Sec. III. The paper then concludes - in Sec. IV - with a discussion of further possible developments.
II Problem Setting of This Paper
As this manuscript bridges the gap between the physics-based assumption model of previous accreditation papers and the cryptographic setting that is more typical of verification protocols, I take an approach that I hope will satisfy both tribes: below I present the problem setting - where Alice, Bob, and their respective objectives are introduced - the crux of which is an protocol (Protocol 1) for how circuits are executed (which Alice and Bob both participate in), and then the physics of errors occurring in a circuit execution are used to support and justify the problem setting, for those who prefer a physics-based error model.
The foundation of my problem setting [^3] Which can be seen as akin to the error model in previous accreditation protocols. is the assumption that error in any operation (e.g. state preparation, gate implementation, or measurement) may be considered as the ideal/errorless operation followed or preceded by a CPTP map, as in Ref. [23]. This is justified as any map from and to density matrices is a CPTP map. Hence, any erroneous operation must be a CPTP map and so any erroneous map may be written as the errorless/ideal operation either followed or preceded by a CPTP map - as unitaries are CPTP maps.
II.1 Introduction to Alice and Bob
@@
With its foundations established, I now present my problem setting, which is an adaption of the cryptographic setting [22]. In the cryptographic setting, Alice is attempting to get Bob to perform a sampBQP [25, 26, 27] quantum computation [^4] Meaning a computation to obtain a sample from a specified distribution, that can be efficiently performed on a quantum computer. Note that sampBQP includes both BQP and fBQP (the set of functions efficiently computable on a quantum computer). and provide her with the true results, while only having the ability to perform polynomially-bounded classical computation herself. Bob is computationally unbounded so, if he wished, could implement the computation Alice requires but his aim is instead to trick Alice by providing her with the results of an incorrect computation and to have her believe he has acted honestly i.e. that the results she receives are the outputs of the computation she wanted to be performed when they are not.
II.2 Problem Setting: Preliminaries and Definitions
@@
My specific problem setting is slightly modified from the traditional cryptographic setting, mainly through limitations on Bob. These limitations are entirely contained in how the circuits Alice requests be executed - and Bob tries to corrupt - are executed, which is presented in Protocol 1 in Sec. II.3, but Protocol 1 first requires the concept of redaction (introduced in Def. 2) and a related concept (in Def. 3).
@@
A gate in a circuit is said to be redacted if the gate’s position in the circuit - i.e. when it is applied and to which qubits - is specified but the operator it represents is not. For an example of a circuit with a redacted gate, see Fig. 1.
Figure 1: An example circuit with a redacted gate (the block in the middle of the circuit). Note how the location of the gate - in terms of time and which qubits are affected - is depicted but which operation the gate represents is hidden.
@@
If given a circuit with all single-qubit gates redacted, the set of all circuits - with no redactions - that the given circuit (with redactions) could be (if the redactions were removed) is called its redaction class. For an example of circuits in the same redaction class, see Fig. 2.
Figure 2: Example circuits within the same redaction class, where the redaction class corresponds to the redacted circuit in Fig. 1. Note that the circuit in Fig. 1 could be any of these if the redaction were removed.
The second and final round of terminology required for the presentation of Protocol 1 is related to how I specify the CPTP maps that are modelling the error in a single execution of a circuit. This is performed via CPTP lists and they are defined in Def. 4.
Definition 4.
A CPTP list is an ordered set of CPTP maps.
However, a CPTP list is just a list of CPTP maps. In order to to be able to describe / determine the error in a given circuit, the CPTP list must contain exactly the right number of CPTP maps. In this case the CPTP list is said to fit the circuit.
Definition 5.
A CPTP list fits a circuit if it may be used to describe error  [^5] Any resulting error - including no error - is valid. It need not be in any way physically justified or correspond to a particular quantum computer’s typical error. in an execution of that circuit.
A CPTP list achieves this by containing exactly one CPTP map corresponding to each location in the circuit where error may occur (These locations are: immediately after gates, immediately after state prep, and immediately before measurement)  [^6] The mapping between the CPTP maps in the CPTP list and the locations error occurs in the circuit is made using the ordering of the CPTP list and matching this with an ordering applied to the locations in the circuit where error may occur..
A CPTP list will fit many different circuits and for any execution of a circuit there exists a CPTP list that both fits the circuit and accurately represents the error occurring in the circuit. I refer to this as the CPTP list determining the error occurring in a circuit execution.
For an example of a CPTP list determining the error in a circuit execution, see Fig. 3.
Figure 3: If $\Xi_{1}=[\xi_{1},\xi_{2},\xi_{3}]$ is a CPTP list used to determine the error in an execution of the top circuit in Fig. 2, then Fig. 3 depicts the circuit that is actually executed. $\xi_{1}$ describes the error due to state preparation, $\xi_{2}$ describes the error due to the single-qubit gate, and $\xi_{3}$ describes the error due to measurement.
Note that notation is slightly abused to display CPTP maps as gates (depicting error) in a circuit, to remedy this slightly they are highlighted in red to denote they are not unitaries but are CPTP maps.
Note 1.
For any set of circuits within the same redaction class, if a given CPTP list fits one of them, it fits all of them.
An important feature to note about CPTP lists and the circuits it fits is mentioned - without proof - in Lemma 1. Proof is omitted as Lemma 1 follows from the above discussion.
Lemma 1.
The error in an execution of a circuit is entirely and exclusively determined by the CPTP list used to describe its error (and the outcomes of any stochastic processes that may occur in the application of the error described by the CPTP list  [^7] As a CPTP map may describe the stochastic application of operations.).
Definition 6.
The probability of error in the execution of a circuit, due to a CPTP list, is the probability that the execution does not provide a sample from the same distribution that executing the circuit as intended (i.e. with no error applied) would. Due to Lemma 1, the probability of error is entirely determined by the CPTP list used to represent the error in the execution. For example, using $\Xi_{1}=[\xi_{1},\xi_{2},\xi_{3}]$ from Fig. 3, if, for all density matrices, $\rho$:
$\xi_{1}(\rho)$ $=\dfrac{1}{2}\rho+\dfrac{1}{2}\hat{X}\rho\hat{X}$ (1)
$\xi_{2}(\rho)$ $=\dfrac{1-0.01}{2}\rho+\dfrac{1+0.01}{2}\hat{X}\rho\hat{X}$ (2)
$\xi_{3}(\rho)$ $=\dfrac{1+0.01}{2}\rho+\dfrac{1-0.01}{2}\hat{X}\rho\hat{X},$ (3)
then the probability of error in the execution of a circuit where these CPTP maps (i.e. $\Xi_{1}$) determine the error is the probability that any of $\xi_{1}$, $\xi_{2}$, or $\xi_{3}$ cause error. In this case, the probability of error is $\approx 0.875$. This quantity is not always meaningful if the error in an execution is not stochastic Pauli error, therefore, for all the traps and target circuits used throughout this paper, the probability of error is defined to be the probability error occurs once the error has been effectively reduced to - and can be considered as - stochastic Pauli error (by the Pauli twirling, defined in Def. 8, and present in all traps and targets considered herein).
Definition 7.
A Set of Probabilistically Similar CPTP Lists with parameter $\beta\in\mathbb{R}$ (henceforth denoted as a PSCLβ) is a CPTP list such that for any CPTP list in the set, the probability of error in any circuit execution it determines the error for is within the interval:
$\big{[}P_{0}\big{(}1-\beta\big{)},P_{0}\big{(}1+\beta\big{)}\big{]},$ (4)
for some $P_{0}\in[0,1]$. Additionally, all CPTP maps in the PSCLβ must be of the same size, so if one fits a given circuit all CPTP maps in the PSCLβ do.
For example consider two CPTP lists: $\Xi_{2}=[\xi_{1}]$ and $\Xi_{3}=[\xi_{2}]$ (where $\xi_{1}$ and $\xi_{2}$ are as in Def. 6). Then the set $\{\Xi_{2},\Xi_{3}\}$ can be considered as a PSCLβ for any $\beta>0.01$ (where the value of $P_{0}$ in this PSCLβ is $0.5$).
With both PSCLβ and redaction classes defined, they may be related via the below Lemma 2. As will be seen in Protocol 1, Alice chooses a redaction class (implicitly, by choosing a subset of one to execute) and Bob chooses a PSCLβ to determine the error in those executions. Hence, PSCLβ and redaction classes are - in some sense - duals of each other. This duality is further expounded apon by Lemma 2.
Lemma 2.
For any given redaction class and any given PSCLβ either:
$\mathbb{1.}$ Every CPTP list in the PSCLβ fits every circuit in the redaction class
$\mathbb{or}$
$\mathbb{2.}$ No CPTP list in the PSCLβ fits any circuit in the redaction class.
I refer to the first of these possibilities as the redaction class and the PSCLβ matching, but also allow the term to apply if the redaction class is replaced by any subset of itself.
Proof.
As in Note 1, if a given CPTP list fits one circuit in a redaction class, it fits every circuit in that redaction class.
As mentioned in Def. 7 if one CPTP list in a given PSCLβ fits a specific circuit then the entire PSCLβ fits that circuit.
The lemma follows from combining these two facts. ∎
II.3 Problem Setting Presentation
@@
I can now present the protocol by which sets of circuits, in my problem setting, are executed, in Protocol 1.
@@
    Alice provides a set of circuits to execute, $\mathcal{S}$, all within the same redaction class.
@@
    Bob receives the full details of $\mathcal{S}$, with all single-qubit gates in every element of $\mathcal{S}$ redacted.
@@
    Bob chooses a SPSCLβ that matches $\mathcal{S}$, denoted $\Xi$, where $\beta\in\mathbb{R}^{+}$ is known to both Alice and Bob.
  4. 4.
    For each circuit, $\mathcal{C}$, in $\mathcal{S}$:
@@
      Bob chooses a CPTP list, $\xi$, from $\Xi$.
@@
      $\mathcal{C}$ is executed with $\xi$ determining the error occurring during the execution.
@@
      Alice and Bob are both given the measurement outcomes of the execution of $\mathcal{C}$.
@@
Alice is free to initiate Protocol 1 whenever and as often as she likes; Bob is free to act how he likes to achieve his stated malicious aims but must execute Protocol 1 as prescribed - but is free to act nefariously within the permitted bounds - when it is invoked. While the bulk of this paper focuses on constructing the accreditation protocol protocol, without much regard for Alice, it culminates in Theorem 2 showing that Alice’s problem is solved by the protocol presented herein.
II.4 Physical Origins of the Error Model / Problem Setting
The two main limitations on the adversarial error / Bob requiring justification are:
@@ II.4 Physical Origins of the Error Model / Problem Setting
    Within a single use of Protocol 1, Bob is restricted to only causing error describable by CPTP lists with a single SPSCLβ, for a known $\beta\in\mathbb{R}^{+}$.
The first limitation corresponds to single-qubit gates experiencing gate-independent error. This is a standard assumption the the pre-existing accreditation protocols (in Refs. [22, 23, 24]) and follows from single-qubit gates typically being the least error-prone components of a quantum computer [28, 29] (which has held true over time [30, 31], and can be most clearly seen in Ref [32, Fig.5]): the error is so small that the error in different single-qubit gates does not differ much. This is a “standard [assumption] in the literature on noise characterisation and mitigation” [33] and has seen extensive use in theoretical work [3, 34, 35, 36, 37, 38, 39, 40, 41, 42, 33].
@@
    Ref. [23, Figure. 5] ran the accreditation protocol - developed therein - many times, producing a probability of error in the traps for each use of the protocol. The set of error probabilities generated by this can be seen to vary little across many uses of the protocol.
@@
    Ref. [43, Fig. 6, Fig. 7, and Fig. 8] investigated the error rates in single-qubit and two-qubit gates. It found these error rates are very rarely far from their average. Note that the variation shown in this paper is over a much longer time-span (days) than this limitation requires (seconds).
@@
    Ref. [32, Fig. 9] examined NISQ devices and - across many devices - plotted the error rates of different qubits in each device. Fig. 9 shows the error bars on the error rate to be - with some notable exceptions - small fractions of the error rate and varying much more across qubits than for a fixed qubit over time.
@@
    Ref. [44, Fig. 1(b)] looked at the error rates of two-qubit gates on differing pairs of qubits, it shows that - again, over a time-span of days - for a specific pair of qubits, the days when the error rate deviates far from its average are rare, with more extreme deviations being rarer.
Finally, this limitation can be seen as an aim of quantum computer hardware engineering: as quantum computers improve and their actual outputs approach the errorless outputs, with decreasing variance, the typical difference - by any measure - between the error contained in two executions of similar circuits will tend to zero.
To my knowledge, I am the first to explicitly state and use this second limitation. However, I note that it is already implicitly accepted in the community by the acceptance of randomized benchmarking [3, 4, 5, 6] as a meaningful measure: if even the same circuit repeated multiple times, in very quick succession, produces wildly varying probabilities of error then the variation of error probabilities implies that previous measures mean almost nothing for future computations or the quality of a quantum device over a meaningful time-span. This is not the case and Ref. [44, Fig. 6] shows that, over a span of days, the results of randomized benchmarking do not vary over a very large range.
There is, however, another, subtle, restriction on Bob: he controls the CPTP maps acting on all executions, but does not control any stochastic elements of the implementation of the CPTP map. I.e. he may specify that a specific CPTP map in the execution applies stochastic Pauli error, such as a dephasing channel, but he does not decide if this results in a Pauli $\hat{Z}$ gate being applied. Furthermore, Bob is not informed of the outcomes of these stochastic processes - in spite of choosing which stochastic processes occur. This represents that the environment - where any memory of past error that the noise represented by Bob has must be - is not, consistently or over long time-spans, protected from noise acting on it and destroying any information/memory it contains, so any hostile intelligence that lurks in the noise must have a comparatively short and small memory  [^8] Quantum hardware engineers have expended much effort to keep the the information inside quantum computers intact, so there is no reason to assume the environment can - with non-negligible probability - store significant amounts of information, for relevant timespans, without any engineering effort to enable it to do so..